# Find Any Part in 3D

### v1

@@ 5.2.2 In-the-Wild Results
Flexibility of text queries. Find3D supports various query types that might occur in-the-wild. As shown in Fig. 7, Find3D can locate hands via different query types – either by the body part “hand” or by the clothing “gloves”. The teddy bear example demonstrates flexibility in query granularity – one can query with “limbs”, a combination of arms and legs, or with “arms” and “legs” separately. For ease of visualization, the scores are min-clipped at 0.
@@ 5.2.3 Failure Modes
We observe some limitations of Find3D: 1) Our model voxel-samples point clouds at the 0.02 resolution (after normalization). Fine-grained parts that are not geometrically prominent, such as bottons on a surface, are difficult for a point-cloud-only model like ours. 2) Because the model is trained to be rotational-equivariant, it tends to make symmetric predictions where all symmetric parts have the same label. Fig. 8 demonstrates an example from the PartNet-E dataset. These limitations point to the complementary nature of the 2D and 3D modalities. While lacking in 3D geometry, the 2D modalities can better convey detailed appearance, which allows us to recognize entities such as buttons or inscriptions on a surface. Combining the image and the point cloud modality is a future direction.
@@ 6 Discussions and Conclusions
Find3D is a zero-shot, open-world, direct-prediction model for 3D part segmentation on any object. It is enabled by a scalable data engine and a contrastive objective that allows for training on large-scale, diverse 3D assets from the internet. Find3D is performant, robust, and efficient. It can work on in-the-wild 3D constructions from iPhone photos or AI-generated images.
We propose some future directions to further improve open-world 3D segmentation. One approach is to combine 2D and 3D modalities to better capture parts without prominent color or geometry characteristics, which are more easily perceived in 2D. Another direction is to investigate the scaling laws of open-world 3D segmentation. Find3D, at a scale of 30K annotated objects and 46.2M parameters, already demonstrates the power of scaling up beyond existing part segmentation datasets. Although its current scale is constrained by budget, we are optimistic that further upscaling of training, and an investigation into the scaling laws, would be beneficial to the community.
@@ 7 Acknowledgments
Seen Categories Unseen Categories
{part} of a {object} {part} {part} of a {object} {part}
PointCLIPV2 9.811 11.267 10.270 11.092
PartSLIP++ 2.690 15.031 0.572 10.427
OpenMask3D 11.810 11.934 7.013 10.307
Find3D (ours) 33.775 34.095 26.205 27.409
Table 3: Detailed results on Objaverse. Two query prompts are evaluated: “{part} of a {object}”, and “{part}”. Our method has a significant advantage in all evaluation configurations.
@@
Ours 43.613 28.386 24.085 43.781 29.637 23.712 50.002 42.151 30.018
Table 4: Detailed results on ShapeNet-Part subset. PointCLIPV2 is trained on this dataset, whereas other methods are evaluated zero-shot. We evaluate different orientations, query prompts, and data domains (ShapeNet-Part vs. Objaverse-ShapeNetPart). We evaluate on 3 types of prompts: “{part} of a {object}”, “{part}”, and top-k. Top-k prompt reproduces the PointCLIPV2 paper, which runs an iterative search over $1400\times n_{\text{parts}}$ prompts per object category to choose the best query text prompts. For fairness of comparison, we follow the same procedure to get top-k prompt metrics, although our method is not designed with prompt searching in mind, and it is not realistic to conduct this hours-long searching process at inference time. Our method, despite being zero-shot on this dataset, has the best performance in 8 out of 9 configurations—all configurations except for the canonical orientation with top-k prompt searching.
@@
Ours 41.517 28.532 23.569 42.734 29.966 23.794
Table 5: Detailed results on ShapeNet-Part full test set. PartSLIP2 and OpenMask3D are too slow and thus infeasible to evaluate on the full test set. The metrics are very close to the subset results in the previous table. Our method, despite being zero-shot on this dataset, has the best performance in 5 out of 6 configurations—all configurations except for the canonical orientation with top-k prompt searching. This searching process takes over an hour on an A100 and our method is not designed for test-time prompt searching.
@@
Ours 17.143 16.211 16.861 16.384 17.703 16.819 17.620 17.164
Table 6: Detailed results on PartNet-E test set. We evaluate different orientations and query prompts. PartSLIP++ is trained on this dataset, and other methods are evaluated zero-shot. Cells with “-” denote that the method is too slow to be evaluated on the full test set. We evaluate with 2 types of prompts: “{part} of a {object}” and “{part}”. PartSLIP++ achieves the highest performance with the “{part}” prompts, yet the performance drops $84\%$ when we vary the query prompt. This dataset is more challenging for our method due to the sparsity of labels and the presence of small parts that are not geometrically or colorfully prominent (e.g., buttons on a surface with the same color). Nevertheless, our method is more robust to rotation and prompt variation, and clearly outperforms the other baselines not trained on this dataset.
A Data Engine
Fig. 9 shows the prompt we use to obtain object orientations from Gemini. For a given orientation, we render the object in 10 different views, and pass the prompt along with 10 renderings to Gemini. We calculate the percentage of “yes” answers and choose the orientation with the highest “yes” percentage. Fig. 9 also provides some example objects with answers from Gemini.
@@ A Data Engine
Fig. 10 shows the prompt we use to obtain part names from Gemini, along with some examples.
@@
B Experiments
B.1 Full Results
In Table 2 of the main paper, in order to evaluate all methods on the exact same data, we had to report results on subsets of ShapeNet-Part and PartNet-E because methods like PartSLIP++ and OpenMask3D are slow and infeasible to evaluate on the full test sets (e.g., OpenMask3D would take 628 hours on PartNet-E). Here we provide full-set results for methods that are feasible for full-set evaluation in Tab. 5 and Tab. 6. The ranking of methods on the full sets and the subsets are the same. Additionally, in Table 2 of the main paper we evaluate all methods with the “part of a object” prompt. We report results evaluated with various prompts on all datasets in Tab. 5, Tab. 5, Tab. 5, and Tab. 6.
The subset indices for ShapeNet-Part can be found at code/Find3D/model/evaluation/benchmark/benchmark_reproducibility/shapenetpart/subset_idxs.json in the supplementary code folder, and indices for PartNet-E can be found at code/Find3D/model/evaluation/benchmark/benchmark_reproducibility/shapenetpart/subsetidxs.json in the supplementary code folder. The random rotations used for evaluation are saved at the same folders. Below we detail results on each dataset.
Objaverse-General. Tab. 5 shows full results on the Objaverse benchmark. Find3D shows significant advantage in all configurations, up to 3x mIoU of the second best method.
ShapeNet-Part. Tab. 5 and Tab. 5 show full results on the ShapeNetPart categories. Tab. 5 compares all methods with various prompts, orientations, and data sources (ShapeNet-Part vs. Objaverse-ShapeNetPart, a benchmark of the same object classes as ShapeNet-Part but sourced from Objaverse that we constructed). We evaluate “{part} of a {object}”, “{part}”, and top-k prompt. The top-k prompt reproduces PointCLIPV2’s evaluation setup. For each object class, PointCLIPV2 runs an iterative search over $1400\times n_{\text{parts}}$ prompts per object category on the test set to choose the best query text prompts. This hours-long test-time search is unrealistic to perform in real applications. Although our method is not designed with prompt searching in mind, we follow the same procedure for our method for fairness of evaluation. PointCLIPV2 is trained on this dataset, and other methods are evaluated zero-shot. Tab. 5 shows that our method performs the best in 8 out of 9 configurations. While Tab. 5 reports metrics on the subset of ShapeNet-Part so that all methods can be evaluated strictly on the same dataset, for methods that are fast enough to evaluate on the full test set (Find3D and PointCLIPV2), we also report the full-set evaluation results in Tab. 5. We note the full-set metrics are very close to the subset metrics. On the full set, we also see our method to perform better in 5 out of 6 settings.
On both the full set and the subset, Find3D, despite being zero-shot on this dataset, is the best-performing method in all configurations except for one—the canonical orientation with test-time top-k prompt searching. In this setting, PointCLIPV2, a method trained on this dataset and designed with test-time prompt searching in mind, performs slightly better. We note that this searching takes over an hour on an A100, which is unrealistic to perform in real applications. Our method is not designed for test-time prompt searching.
PartNet-E. Tab. 6 provides comprehensive evaluations on PartNet-E, both on the subset (for all methods) and on the full set (for methods that are fast enough to evaluate on the full set). PartSLIP++, trained on this dataset, achieves the highest performance with the “{part}” prompts, yet is very sensitive to prompt variation. We note that PartSLIP++ also releases category-specific checkpoints, but we use the cross-category checkpoint for fairness of comparison. This dataset is more challenging for our method because 1) many objects contain small parts that are not geometrically or colorfully prominent, such as buttons on a surface with the same color. 2) Some objects have very sparse labels, such as a bottle with only less than $10\%$ of points labeled as “lid”. Our method has a tendency to predict densely (i.e. less likely to predict “no label”), and thus performs less well on this dataset. Nevertheless, we see our method to be more robust to rotation and prompt variation, and clearly outperforms the other baselines that are not trained on this dataset. Furthermore, PartSLIP++ is a slow 2D-3D aggregation method, taking up to 3 minutes per object. Our method is over 30$\times$ faster.

### v2

@@ 5.2.3 Flexibility of text queries
Find3D supports various query types that might occur in-the-wild. As shown in Fig. 8, Find3D can locate hands via different query types – either by the body part “hand” or by the clothing “gloves”. The teddy bear example demonstrates flexibility in query granularity – one can query with “limbs”, a combination of arms and legs, or with “arms” and “legs” separately. For ease of visualization, the scores are min-clipped at 0.
@@ 5.2.4 Failure Modes
We observe some limitations of Find3D: 1) Our model voxel-samples point clouds at the 0.02 resolution (after normalization). Fine-grained parts that are not geometrically prominent, such as bottons on a surface, are difficult for a point-cloud-only model like ours. 2) Because the model is trained to be rotational-equivariant, it tends to make symmetric predictions where all symmetric parts have the same label. Fig. 9 demonstrates an example from the PartNet-E dataset. These limitations point to the complementary nature of the 2D and 3D modalities. While lacking in 3D geometry, the 2D modalities can better convey detailed appearance. Combining the image and the point cloud modality is a future direction.
@@ 6 Discussions and Conclusions
We present the first scaling study for 3D part segmentation. Key to our approach is a data engine that automatically annotates 3D assets from the internet, which allows us to train the first zero-shot generalist model for open-world 3D part segmentation on any object. Our method not only shows strong generalization, but even outperforms prior methods on the datasets they train on, despite being zero-shot. We show that training object diversity is critical with a scaling analysis. We will release our code, benchmark and model checkpoints. We hope that by providing a diverse benchmark and the first demonstration of open-world 3D part segmentation at scale, we can encourage the community to shift away from customizations for small-scale datasets towards scale and generalization.
@@ 7 Acknowledgments
A Additional data engine annotation examples
We provide additional examples of our data engine annotations, both for high-quality examples in  Fig. 10 and lower-quality (but still useful) examples in  Fig. 11. Upon manual inspection of 50 randomly sampled objects, we observe $76\%$ high-quality examples. The annotations cover diverse objects and descriptions. For example, the body of a fire extinguisher is referred to both as “body” and “cylinder” from different views. The lower-quality examples are still useful for training – they might not have pronounced parts, or contain partial masks (e.g., the baguette example), but the supervision signal still pushes the point features close to the correct semantic embedding (e.g. bread-related concepts). The low-quality cartoon frog contains both correct and incorrect masks. When learning from millions of such labels, the incorrect labels can be “smoothed out” because it’s unlikely that many frogs’ bellies are all incorrectly labeled as “bowtie”.
A Additional qualitative examples of Find3D
We provide additional qualitative results of Find3D in  Fig. 12 and  Fig. 13.  Fig. 12 shows predictions on Objaverse-General from 4 views for each object.  Fig. 13 shows predictions on PartObjaverse-Tiny [yang2024sampart3d] and iPhone photos (reconstructed to 3D via off-the-shelf single-image reconstruction method, Trellis [25])). Find3D can segment diverse objects and parts, and can generalize to real-world objects, despite being trained on synthetic data.
A Experiments
A.1 Additional Results
In Tab. 2, Tab. 3, Tab. 4 of the main paper, in order to evaluate all methods on the exact same data, we had to report results on subsets of ShapeNet-Part and PartNet-E because methods like PartSLIP++ and OpenMask3D are slow and infeasible to evaluate on the full test sets (e.g., OpenMask3D would take 628 hours on PartNet-E). Here we provide full-set results for methods that are feasible for full-set evaluation in Tab. 6 and Tab. 7. The ranking of methods on the full sets and the subsets are the same.
ShapeNet-Part. Tab. 5 compares all methods with various prompts, orientations, and data sources (ShapeNet-Part vs. ShapeNetPart-V2, a benchmark of the same object classes as ShapeNet-Part but sourced from Objaverse that we constructed, similar to ImageNetV2 [16]). PointCLIPV2 is trained on this dataset, and other methods are evaluated zero-shot. Find3D performs the best in 8 out of 9 configurations, despite being zero-shot. While Tab. 5 reports metrics on the subset of ShapeNet-Part so that all methods can be evaluated strictly on the same dataset, for methods that are fast enough to evaluate on the full test set (Find3D and PointCLIPV2), we also report the full-set evaluation results in Tab. 6. The full-set metrics are very close to the subset metrics. On the full set, we also see that Find3D performs better in 5 out of 6 settings.
On both the full set and the subset, Find3D, despite being zero-shot on this dataset, is the best-performing method in all configurations except for one—the canonical orientation with test-time top-k prompt searching. In this setting, PointCLIPV2, a method trained on this dataset and designed with test-time prompt searching in mind, performs slightly better. We note that this searching takes over an hour on an A100, which is unrealistic to perform in real applications. Our method is not designed for test-time prompt searching but clearly outperforms all baselines when doing direct inference.
[…]
A Data engine prompts
Fig. 14 shows the prompt we use to obtain object orientations from Gemini. For a given orientation, we render the object in 10 different views, and pass the prompt along with 10 renderings to Gemini. We calculate the percentage of “yes” answers and choose the orientation with the highest “yes” percentage. Fig. 14 also provides some example objects with answers from Gemini. Fig. 15 shows the prompt we use to obtain part names from Gemini, along with some examples.
Figure 10: High-quality examples of data engine annotations. The LVIS label (from Objaverse) is shown below each input object. Our data engine annotates diverse objects and parts, including multiple captions for the same parts, such as “candelabra arm” and “candlestick arm”, and multiple levels of granularity, such as “helmet shell” and “ear pad”.
Figure 11: Lower-quality examples of data engine annotations. The LVIS label (from Objaverse) is shown below each input object. Some objects do not have pronounced parts, such as the baguette, and get partial part labels due to texture/lighting change on surfaces. Some objects are low quality, such as the cartoon frog, which results in incorrect labels.
Figure 12: Multiple views of Find3D predictions on Objaverse-General examples.
Figure 13: Multiple views of Find3D predictions on PartObjaverse-Tiny examples and iPhone photos (reconstructed to 3D with off-the-shelf method).
@@ A Data engine prompts
Find3D (Ours) 43.613 28.386 24.085 43.781 29.637 23.712 50.002 42.151 30.018
Table 5: Detailed results on ShapeNet-Part subset. Shaded cells mean the method is trained on the same dataset (expected higher than white cells), and white cells mean zero-shot evaluation. We evaluate different orientations, query prompts, and data domains (ShapeNet-Part vs. ShapeNetPart-V2). We evaluate on 3 types of prompts: “{part} of a {object}”, “{part}”, and top-k. Top-k prompt reproduces the PointCLIPV2 paper, which runs an iterative search over $1400\times n_{\text{parts}}$ prompts per object category to choose the best query text prompts. For fairness of comparison, we follow the same procedure to get top-k prompt metrics, although our method is not designed with prompt searching in mind, and it is not realistic to conduct this slow ($>1$ hour on A100) searching process at inference time. Our method, despite being zero-shot on this dataset, has the best performance in 8 out of 9 configurations—all configurations except for the canonical orientation with top-k prompt searching.
@@
Find3D (Ours) 41.517 28.532 23.569 42.734 29.966 23.794
Table 6: Detailed results on ShapeNet-Part full test set. Shaded cells mean the method is trained on the same dataset (expected higher than white cells), and white cells mean zero-shot evaluation. PartSLIP2 and OpenMask3D are too slow and thus infeasible to evaluate on the full test set. The metrics are very close to the subset results in the previous table. Our method, despite being zero-shot on this dataset, has the best performance in 5 out of 6 configurations—all configurations except for the canonical orientation with top-k prompt searching. This searching process takes over an hour on an A100 and our method is not designed for test-time prompt searching.
@@
Find3D (Ours) 17.143 16.211 16.861 16.384 17.703 16.819 17.620 17.164
Table 7: Detailed results on PartNet-E test set. Shaded cells mean the method is trained on the same dataset (expected higher than white cells), and white cells mean zero-shot evaluation. Cells with “-” denote that the method is too slow to be evaluated on the full test set. We evaluate with 2 types of prompts: “{part} of a {object}” and “{part}”. PartSLIP++ achieves the highest performance with the “{part}” prompts, yet the performance drops $84\%$ when we vary the query prompt. This dataset is more challenging for our method due to the sparsity of labels and the presence of small parts that are not geometrically or colorfully prominent (e.g., buttons on a surface with the same color). Nevertheless, our method is more robust to rotation and prompt variation, and clearly outperforms the other baselines not trained on this dataset.